OpenStudy (anonymous):

Kathy & Leslie drove total of 225 miles in 5.2 hours. Kathy drove first part of the trip & averaged 45 miles per hour. Leslie drove remainder & averaged 55 miles per hour. For what length of time did Kathy drive?

OpenStudy (anonymous):

first part as in like first half?

OpenStudy (anonymous):

Yes.

OpenStudy (pfenn1):

If x = time Kathy drove and y=time Leslie drove x+y=5.2 hours and 45x + 55y = 225 miles

OpenStudy (anonymous):

Let us assume Kathy Drove for x hours so Leslie drove for (5.2-x) hours 45x + 55(5.2-x) = 225 solve for x

OpenStudy (anonymous):

@sheg 45+286-55x=225?

OpenStudy (anonymous):

@pfenn1 Thank you for your help too.

OpenStudy (pfenn1):

you are welcome.

OpenStudy (anonymous):

So, I'm kind of stuck on what @sheg said. Is it 45+286-55x=225?

OpenStudy (pfenn1):

yes, now solve for x

OpenStudy (anonymous):

I got like 1.9.

OpenStudy (anonymous):

The answer is 3.1 hours but I don't understand how to get that.

OpenStudy (anonymous):

wait i feel we are going in the wrong direction

OpenStudy (anonymous):

Ah, me too.

OpenStudy (anonymous):

Im sorry this troubling.

OpenStudy (pfenn1):

The problem is when you expanded the equation 45x + 55(5.2-x) = 225 should be 45x+286-55x=225 I didn't notice that the x had been dropped from the first term. Try it again now that the equation has been corrected.

OpenStudy (anonymous):

Your right! Okay, let me try this again.

OpenStudy (pfenn1):

@sheg, you developed it correctly, we just didn't notice the dropped x.

OpenStudy (anonymous):

So I did the problem and I got the answer 6.1.

OpenStudy (pfenn1):

check it again.....you should have gotten 3.1....

OpenStudy (anonymous):

Ugh okay lol.

OpenStudy (anonymous):

cause its 45x-55x=-10x+286 then subtract equals -61 so then -10x=-61 then divide i got 6.1?

OpenStudy (anonymous):

your question i feel is not correct how can they drive @ a speed of 45 and 55 but still take 5.2hours i m getting 4.3 hours

OpenStudy (anonymous):

I could upload the picture.

OpenStudy (pfenn1):

Did you copy the total distance correctly in the original problem. I'm guessing it was 255 miles total, not 225 miles. It is getting really late and I think we are all getting tired and making mistakes. I was using 255 even though it said 225 so I got the right answer when I did it.

OpenStudy (anonymous):

hey average speed of kathy and leslie is (45+55)/2 = 50miles/hr and they can cover 225miles in 225/50 = 9/2 = 4.5 hours

OpenStudy (anonymous):

Im so dumb it was 255 :(

OpenStudy (anonymous):

Im really sorry guys. :( I got it from here now. Thank you.

OpenStudy (anonymous):

my god @amandamv1994 you r great

OpenStudy (pfenn1):

No, you are not dumb. It was just a typo that we all missed.

OpenStudy (anonymous):

Haha i can't believe I read it wrong! Grrrr.

OpenStudy (anonymous):

But your help still helped me so thank you!

OpenStudy (pfenn1):

I can't believe I read it wrong and got the wrong answer using the the wrong "tight" number LOL!

OpenStudy (pfenn1):

I meant got the *right answer and I meant "right" not "tight" I'm going to bed. I am too old to be so late.

OpenStudy (pfenn1):

I meant up so late

OpenStudy (anonymous):

Sorry to both of you :(

OpenStudy (pfenn1):

no problem. sorry it took so long for us to find the mistakes.....you were doing it right all along but just with typos. good night.

OpenStudy (anonymous):

Goodnight!